OpenStudy (r_josh10):

Robby invested $18,400 at 7% interest. He earned $9016 interest at the end of x years. What is the value of x? A. 10 B. 9 C. 8 D. 7 I will fan and medal.

OpenStudy (nurali):

you would have to use the formula P*T*R=I P=Principle=18400 R=Rate=7%=0.07 T=The number of years=X I=Interest=9016

OpenStudy (r_josh10):

So what would I do?

OpenStudy (nurali):

X=9016/18400*0.07 X=9016/1288

OpenStudy (r_josh10):

Thank You

OpenStudy (nurali):

My Pleasure.